# CompEvent: Complex-valued Event-RGB Fusion for Low-light Video Enhancement and Deblurring

### v1

@@ Abstract
Low-light video deblurring poses significant challenges in applications like nighttime surveillance and autonomous driving due to dim lighting and long exposures. While event cameras offer potential solutions with superior low-light sensitivity and high temporal resolution, existing fusion methods typically employ staged strategies, limiting their effectiveness against combined low-light and motion blur degradations. To overcome this, we propose CompEvent, a complex neural network framework enabling holistic full-process fusion of event data and RGB frames for enhanced joint restoration. CompEvent features two core components: 1) Complex Temporal Alignment GRU, which utilizes complex-valued convolutions and processes video and event streams iteratively via GRU to achieve temporal alignment and continuous fusion; and 2) Complex Space-Frequency Learning module, which performs unified complex-valued signal processing in both spatial and frequency domains, facilitating deep fusion through spatial structures and system-level characteristics. By leveraging the holistic representation capability of complex-valued neural networks, CompEvent achieves full-process spatiotemporal fusion, maximizes complementary learning between modalities, and significantly strengthens low-light video deblurring capability. Extensive experiments demonstrate that CompEvent outperforms SOTA methods in addressing this challenging task. The code is available at https://github.com/YuXie1/CompEvent.
@@ Introduction
In applications such as nighttime surveillance and autonomous driving, video capture in low-light environments inevitably requires extended exposure times, often suffering from the dual degradations of insufficient brightness and motion blur, leading to a sharp decline in video quality (kurimo2009effect; kim2024towards). These two degradations are tightly coupled: the long exposure required to increase brightness actually exacerbates the blur of moving objects and obliterates significant edge and texture details. This makes the joint task of video enhancement and deblurring a highly ill-posed problem.
@@
Traditional video restoration methods primarily rely on information within the frame itself and face a fundamental information bottleneck when dealing with this compound degradation. Early model-driven approaches (horn1981determining; lucas1981iterative) and more recent deep learning-driven approaches, including convolutional neural networks (CNNs) (nah2017deep) and Transformers (vaswani2017attention; wang2022uformer; zamir2022restormer), perform poorly in low-light scenarios. This is because, in low-light conditions, the extended exposure time required to compensate for brightness exacerbates motion blur and significantly loses edge and texture details. Compared to deblurring under normal exposure, models face a more challenging task in estimating the blurring process in low-light conditions. Simply combining low-light enhancement and deblurring tasks in tandem often leads to suboptimal results due to the accumulation and amplification of errors (e.g., noise is enhanced and blur is solidified) (zhou2022lednet).
@@
Event cameras, with their unique advantages of high temporal resolution (HTR) and high dynamic range (HDR), offer new possibilities for addressing this problem (Ruedi 1996; Lichtsteiner and Delbruck 2005) (lichtsteiner200364x64). Their HTR features capture fine motion trajectories to guide deblurring, while their HDR features perceive scene structure in extremely dark environments to aid low-light enhancement (brandli2014240). The precise spatiotemporal information provided by these two features perfectly complements the rich texture and color of RGB frames.
@@
However, effectively fusing the advantages of these two heterogeneous modalities remains an open challenge. Existing event-RGB fusion methods, including some pioneering work (kim2024towards), mostly follow a ”staged fusion” strategy. In this paradigm, the network processes the two modalities in independent streams, exchanging information only at specific, discrete nodes, as shown in Figure 1. This discontinuous fusion approach fundamentally limits the network’s ability to learn a deep, collaborative feature representation. Between the two fusion nodes, the network is forced to independently learn suboptimal single-modal features, making the fusion itself more of a ”patching” operation than a deep integration. This approach fails to fully exploit the fine-grained spatiotemporal correlations between the two data sets, severely limiting its performance in scenarios that rely heavily on complementary information, such as intense motion or extreme dimming. Therefore, we believe that a more optimal solution should implement deep interaction throughout, allowing the features of the two modalities to co-evolve at every layer of the network processing, thereby building a comprehensive and robust understanding of degraded scenes.
@@
Traditional video deblurring relies on frames alone. Early methods were model-based (levin2009understanding), while modern approaches leverage deep learning networks, from CNNs (nah2017deep) to Transformers (wang2022uformer; zamir2022restormer). However, in low-light scenes, long exposure times worsen motion blur and severely degrade edge and texture detail (kim2024towards), making motion estimation and detail recovery extremely difficult. To address this, researchers use event cameras, which asynchronously record brightness changes with high temporal resolution, capturing motion trajectories lost in blurred frames. Events thus serve as effective motion priors for handling severe blur (qi2024deblurring; liang2023coherent).
@@
Low-light enhancement methods are also predominantly frame-based. They include Retinex-based models (land1977retinex; wei2018deep) that decompose images into illumination and reflection, and zero-reference methods (guo2020zero; jiang2021enlightengan) that learn enhancement without ground truth. However, applying these methods directly to blurry low-light videos amplifies noise and artifacts (zhou2022lednet). Event cameras, due to their high dynamic range, can preserve scene structure even in underexposed areas. Introducing events provides structural priors absent in frames, aiding detail restoration and brightness improvement without blindly enhancing degradation.
@@
Low-light and motion blur are physically coupled, motivating unified solutions. Frame-based methods such as LEDNet (zhou2022lednet) and JUDE (vo2025deep) improve upon cascaded models via joint architectures and deep algorithm but remain limited by frame information (kim2024towards). Kim et al. (2024) introduced the RELED dataset and ED-TFA, a staged event-frame fusion module. Yet staged fusion—common in prior work—restricts cross-modal interaction to specific stages, limiting recovery under extreme degradation. We instead propose a full-process fusion strategy using complex neural networks, enabling continuous spatiotemporal integration to more effectively solve this joint restoration task.
@@
As shown in the Figure 2, complex convolution jointly operates on the real part $M_{R}$ and the imaginary part $M_{E}$ via a shared kernel consisting of $K_{R}$ and $K_{E}$. This operation mechanism allows each complex output $M^{\prime}$ to depend on both modalities, achieving tighter fusion than real convolution. Furthermore, compared to real convolution, this shared structure reduces the number of parameters by nearly $50\%$ while enhancing cross-modal learning. Complex convolution naturally supports a ”full-process fusion” strategy, promoting the continuous interaction of RGB and event features throughout the network.
@@
Accurate temporal alignment of features becomes highly challenging under severe motion blur and low-light noise. Traditional optical flow methods (horn1981determining; lucas1981iterative), rely on the assumptions of constant brightness and spatial smoothness. However, these assumptions often fail under conditions of blur caused by long exposures and noise caused by low light, resulting in motion estimation failures.To address this, we adopt recurrent neural networks (RNNs) to model temporal relations implicitly, avoiding explicit motion estimation. Specifically, we employ gated recurrent units (GRUs) , which flexibly regulates information flow through a gating mechanism and can maintain higher stability in the face of uncertainty and noise (zhou2022lednet).
@@
Where, $\odot$ is the complex Hadamard product. We adopt split activations: for input $z=x+iy$, we define $\sigma_{c}(z)=\sigma(x)+i\sigma(y)$, which has been proven effective and stable in practice (nah2017deep; zamir2022restormer).
@@
To fully utilize the contextual information in a video, we use bidirectional CTA-GRU (wang2022uformer; jiang2021enlightengan). A forward pass processes $t-1\to t\to t+1$, yielding $\{\overrightarrow{H}_{t-1},\overrightarrow{H}_{t},\overrightarrow{H}_{t+1}\}$, while a backward pass processes $t+1\to t\to t-1$, yielding $\{\overleftarrow{H}_{t+1},\overleftarrow{H}_{t},\overleftarrow{H}_{t-1}\}$. The aligned feature at $t$ is:
@@
After temporal alignment through the CTA-GRU module, the features are fed into the backbone network for final image restoration. The mixed degradation of low light and motion blur exhibits different characteristics in different image domains: low light primarily affects low-frequency components (such as overall brightness and contrast) (huang2022deep; chen2023low), while motion blur primarily manifests as an attenuation of high-frequency components (such as edges and texture) (manolakis2011applied). Therefore, an ideal restoration network should be able to synergistically address the fine spatial structure and systematic frequency characteristics of the image.
@@
We use whitening-based CLN (trabelsi2017deep) to preserve correlations between real and imaginary components, ensuring stability during training .
The features are then fed into a Dual Branch Structure (DBS) . The first branch of DBS is the Complex Space Branch, which aims to learn spatially varying, deeply fused representations, extracting and fusing features simultaneously. Its main structure draws on depth-wise separable convolution (chollet2017xception; howard2017mobilenets)—an efficient form of convolution widely used in modern network architectures. We extend it to the complex domain to process the fused features. The operation can be expressed as:
@@
Where, $\text{CFFN}(X)=\text{CConv}_{\text{pw4}}\bigl(\text{sGeLU}\bigl(\text{CConv}_{\text{pw3}}(X)\bigr)\bigr)$, $\text{CConv}_{\text{pw3}}$ and $\text{CConv}_{\text{pw4}}$ are $1\times 1$ complex point-wise convolutions. CFFN performs nonlinear transformations directly in the complex domain. It not only models complex feature relationships like real-valued FFN, but also preserves and utilizes the phase information of the signal, thus possessing richer representation capabilities (bassey2021survey). By stacking multiple such CSFBs, our network can deeply and collaboratively process spatial and frequency domain information at different scales, thereby achieving effective restoration of low-light blurry videos.
@@
Low-Light Enhancement SNRNet (xu2022snr) F 26.47 0.851 20.25 0.815
LLFormer (wang2023ultra) F 26.62 0.862 20.68 0.832
RetinexFormer (cai2023retinexformer) F 26.66 0.865 20.83 0.817
SDSDNet (wang2021seeing) F 28.47 0.887 21.34 0.832
EvLight++ (chen2024evlight++) F+E 30.87 0.888 24.99 0.880
Motion Deblur MPRNet (zamir2021multi) F 26.89 0.867 21.35 0.825
MIMOUNet+ (cho2021rethinking) F 26.52 0.866 21.12 0.821
NAFNet (chen2022simple) F 26.77 0.862 21.28 0.818
RNN-MBP (zhu2022deep) F 29.52 0.902 23.58 0.862
DSTNet (pan2023deep) F 29.59 0.903 23.63 0.864
e-SLNet (wang2020event) F+E 19.45 0.663 17.05 0.738
REDNet (xu2021motion) F+E 29.19 0.903 23.25 0.859
EFNet (sun2022event) F+E 29.85 0.905 23.92 0.867
MAT (xu2025motion) F+E 31.22 0.896 25.15 0.882
UEVD (kim2022event) F+E 29.93 0.905 24.08 0.869
REFID (sun2023event) F+E 30.10 0.913 24.55 0.875
Joint (Frame-based) LEDNet (zhou2022lednet) F 30.36 0.887 25.74 0.850
Joint (Event-guided) ELEDNet (kim2024towards) F+E 31.30 0.925 25.04 0.873
@@
We evaluate CompEvent on the RELED real-world dataset (kim2024towards) and the LOL-Blur synthetic dataset (zhou2022lednet). Training details and hyperparameters are provided in the supplementary material. (1) RELED (kim2024towards) is the first large-scale real-world benchmark built for the joint low-light enhancement and motion deblurring tasks. The dataset is acquired through an optical beam splitting system that can simultaneously record low-light blurry videos, the corresponding high-quality clear images, and high-fidelity event streams. (2) LOL-Blur (zhou2022lednet) is a large-scale synthetic dataset that provides low-light blurry image and clear image pairs for the joint low-light enhancement and deblurring tasks. To adapt our event-based approach, we generate the corresponding event streams for it using the ESIM simulator (rebecq2018esim).

### v3

@@ Abstract
Low-light video deblurring poses significant challenges in applications like nighttime surveillance and autonomous driving due to dim lighting and long exposures. While event cameras offer potential solutions with superior low-light sensitivity and high temporal resolution, existing fusion methods typically employ staged strategies, limiting their effectiveness against combined low-light and motion blur degradations. To overcome this, we propose CompEvent, a complex neural network framework enabling holistic full-process fusion of event data and RGB frames for enhanced joint restoration. CompEvent features two core components: 1) Complex Temporal Alignment GRU, which utilizes complex-valued convolutions and processes video and event streams iteratively via GRU to achieve temporal alignment and continuous fusion; and 2) Complex Space-Frequency Learning module, which performs unified complex-valued signal processing in both spatial and frequency domains, facilitating deep fusion through spatial structures and system-level characteristics. By leveraging the holistic representation capability of complex-valued neural networks, CompEvent achieves full-process spatiotemporal fusion, maximizes complementary learning between modalities, and significantly strengthens low-light video deblurring capability. Extensive experiments demonstrate that CompEvent outperforms SOTA methods in addressing this challenging task.
11footnotetext: Corresponding author.
Codehttps://github.com/YuXie1/CompEvent
@@ Introduction
In applications such as nighttime surveillance and autonomous driving, video capture in low-light environments inevitably requires extended exposure times, often suffering from the dual degradations of insufficient brightness and motion blur, leading to a sharp decline in video quality (Kim et al. 2024). These two degradations are tightly coupled: the long exposure required to increase brightness actually exacerbates the blur of moving objects and obliterates significant edge and texture details. This makes the joint task of video enhancement and deblurring a highly ill-posed problem.
@@
Traditional video restoration methods primarily rely on information within the frame itself and face a fundamental information bottleneck when dealing with this compound degradation. Early model-driven approaches (Horn and Schunck 1981; Lucas and Kanade 1981) and more recent deep learning-driven approaches, including convolutional neural networks (CNNs) (Nah et al. 2017) and Transformers (Vaswani et al. 2017; Wang et al. 2022; Zamir et al. 2022), perform poorly in low-light scenarios. This is because, in low-light conditions, the extended exposure time required to compensate for brightness exacerbates motion blur and significantly loses edge and texture details. Compared to deblurring under normal exposure, models face a more challenging task in estimating the blurring process in low-light conditions. Simply combining low-light enhancement and deblurring tasks in tandem often leads to suboptimal results due to the accumulation and amplification of errors (e.g., noise is enhanced and blur is solidified) (Zhou et al. 2022).
@@
Event cameras, with their unique advantages of high temporal resolution (HTR) and high dynamic range (HDR), offer new possibilities for addressing this problem (Ruedi 1996; Lichtsteiner and Delbruck 2005) (Lichtsteiner 2003). Their HTR features capture fine motion trajectories to guide deblurring, while their HDR features perceive scene structure in extremely dark environments to aid low-light enhancement (Brandli et al. 2014). The precise spatiotemporal information provided by these two features perfectly complements the rich texture and color of RGB frames.
@@
However, effectively fusing the advantages of these two heterogeneous modalities remains an open challenge. Existing event-RGB fusion methods, including some pioneering work (Kim et al. 2024), mostly follow a ”staged fusion” strategy. In this paradigm, the network processes the two modalities in independent streams, exchanging information only at specific, discrete nodes, as shown in Figure 1. This discontinuous fusion approach fundamentally limits the network’s ability to learn a deep, collaborative feature representation. Between the two fusion nodes, the network is forced to independently learn suboptimal single-modal features, making the fusion itself more of a ”patching” operation than a deep integration. This approach fails to fully exploit the fine-grained spatiotemporal correlations between the two data sets, severely limiting its performance in scenarios that rely heavily on complementary information, such as intense motion or extreme dimming. Therefore, we believe that a more optimal solution should implement deep interaction throughout, allowing the features of the two modalities to co-evolve at every layer of the network processing, thereby building a comprehensive and robust understanding of degraded scenes.
@@
Traditional video deblurring relies on frames alone. Early methods were model-based (Levin et al. 2009). With the rapid development of deep learning (Li et al. 2023, 2024a, 2024b; Jiang et al. 2024; Zhu et al. 2024; Li et al. 2025a, b), modern approaches leverage deep learning networks, from CNNs (Nah et al. 2017) to Transformers (Wang et al. 2022; Zamir et al. 2022). However, in low-light scenes, long exposure times worsen motion blur and severely degrade edge and texture detail (Kim et al. 2024), making motion estimation and detail recovery extremely difficult. To address this, researchers use event cameras, which asynchronously record brightness changes with high temporal resolution, capturing motion trajectories lost in blurred frames. Events thus serve as effective motion priors for handling severe blur (Qi et al. 2024; Liang et al. 2023).
@@
Low-light enhancement methods are also predominantly frame-based. They include Retinex-based models (Land 1977; Wei et al. 2018) that decompose images into illumination and reflection, and zero-reference methods (Guo et al. 2020; Jiang et al. 2021) that learn enhancement without ground truth. However, applying these methods directly to blurry low-light videos amplifies noise and artifacts (Zhou et al. 2022). Event cameras, due to their high dynamic range, can preserve scene structure even in underexposed areas. Introducing events provides structural priors absent in frames, aiding detail restoration and brightness improvement without blindly enhancing degradation (Fu et al. 2024; Xu et al. 2024; Liu et al. 2025; Sun et al. 2025).
@@
Low-light and motion blur are physically coupled, motivating unified solutions. Frame-based methods such as LEDNet (Zhou et al. 2022) and JUDE (Vo and Park 2025) improve upon cascaded models via joint architectures and deep algorithm but remain limited by frame information (Kim et al. 2024).  (Kim et al. 2024) introduced the RELED dataset and ED-TFA, a staged event-frame fusion module. Yet staged fusion—common in prior work—restricts cross-modal interaction to specific stages, limiting recovery under extreme degradation. We instead propose a full-process fusion strategy using complex neural networks, enabling continuous spatiotemporal integration to more effectively solve this joint restoration task.
@@
As shown in the Figure 2, complex convolution jointly operates on the real part $M_{R}$ and the imaginary part $M_{E}$ via a shared kernel consisting of $K_{R}$ and $K_{E}$ (Luo et al. 2025). This operation mechanism allows each complex output $M^{\prime}$ to depend on both modalities, achieving tighter fusion than real convolution. Furthermore, compared to real convolution, this shared structure reduces the number of parameters by nearly $50\%$ while enhancing cross-modal learning. Complex convolution naturally supports a ”full-process fusion” strategy, promoting the continuous interaction of RGB and event features throughout the network.
@@
Accurate temporal alignment of features becomes highly challenging under severe motion blur and low-light noise. Traditional optical flow methods (Horn and Schunck 1981; Lucas and Kanade 1981), rely on the assumptions of constant brightness and spatial smoothness. However, these assumptions often fail under conditions of blur caused by long exposures and noise caused by low light, resulting in motion estimation failures.To address this, we adopt recurrent neural networks (RNNs) to model temporal relations implicitly, avoiding explicit motion estimation. Specifically, we employ gated recurrent units (GRUs) , which flexibly regulates information flow through a gating mechanism and can maintain higher stability in the face of uncertainty and noise (Zhou et al. 2022).
@@
Where, $\odot$ is the complex Hadamard product. We adopt split activations: for input $z=x+iy$, we define $\sigma_{c}(z)=\sigma(x)+i\sigma(y)$, which has been proven effective and stable in practice (Nah et al. 2017; Zamir et al. 2022).
@@
To fully utilize the contextual information in a video, we use bidirectional CTA-GRU (Wang et al. 2022; Jiang et al. 2021). A forward pass processes $t-1\to t\to t+1$, yielding $\{\overrightarrow{H}_{t-1},\overrightarrow{H}_{t},\overrightarrow{H}_{t+1}\}$, while a backward pass processes $t+1\to t\to t-1$, yielding $\{\overleftarrow{H}_{t+1},\overleftarrow{H}_{t},\overleftarrow{H}_{t-1}\}$. The aligned feature at $t$ is:
@@
After temporal alignment through the CTA-GRU module, the features are fed into the backbone network for final image restoration. The mixed degradation of low light and motion blur exhibits different characteristics in different image domains: low light primarily affects low-frequency components (such as overall brightness and contrast) (Huang et al. 2022; Chen and Jin 2023), while motion blur primarily manifests as an attenuation of high-frequency components (such as edges and texture) (Manolakis and Ingle 2011). Therefore, an ideal restoration network should be able to synergistically address the fine spatial structure and systematic frequency characteristics of the image.
@@
We use whitening-based CLN (Trabelsi et al. 2017) to preserve correlations between real and imaginary components, ensuring stability during training .
The features are then fed into a Dual Branch Structure (DBS) . The first branch of DBS is the Complex Space Branch, which aims to learn spatially varying, deeply fused representations, extracting and fusing features simultaneously. Its main structure draws on depth-wise separable convolution (Chollet 2017; Howard et al. 2017)—an efficient form of convolution widely used in modern network architectures. We extend it to the complex domain to process the fused features. The operation can be expressed as:
@@
Where, $\text{CFFN}(X)=\text{CConv}_{\text{pw4}}\bigl(\text{sGeLU}\bigl(\text{CConv}_{\text{pw3}}(X)\bigr)\bigr)$, $\text{CConv}_{\text{pw3}}$ and $\text{CConv}_{\text{pw4}}$ are $1\times 1$ complex point-wise convolutions. CFFN performs nonlinear transformations directly in the complex domain. It not only models complex feature relationships like real-valued FFN, but also preserves and utilizes the phase information of the signal, thus possessing richer representation capabilities (Bassey et al. 2021). By stacking multiple such CSFBs, our network can deeply and collaboratively process spatial and frequency domain information at different scales, thereby achieving effective restoration of low-light blurry videos.
@@
Low-Light Enhancement SNRNet (Xu et al. 2022) F 26.47 0.851 20.25 0.815
LLFormer (Wang et al. 2023) F 26.62 0.862 20.68 0.832
RetinexFormer (Cai et al. 2023) F 26.66 0.865 20.83 0.817
SDSDNet (Wang et al. 2021) F 28.47 0.887 21.34 0.832
EvLight++ (Chen et al. 2024) F+E 30.87 0.888 24.99 0.880
Motion Deblur MPRNet (Zamir et al. 2021) F 26.89 0.867 21.35 0.825
MIMOUNet+ (Cho et al. 2021) F 26.52 0.866 21.12 0.821
NAFNet (Chen et al. 2022) F 26.77 0.862 21.28 0.818
RNN-MBP (Zhu et al. 2022) F 29.52 0.902 23.58 0.862
DSTNet (Pan et al. 2023) F 29.59 0.903 23.63 0.864
e-SLNet (Wang et al. 2020) F+E 19.45 0.663 17.05 0.738
REDNet (Xu et al. 2021) F+E 29.19 0.903 23.25 0.859
EFNet (Sun et al. 2022) F+E 29.85 0.905 23.92 0.867
MAT (Xu et al. 2025) F+E 31.22 0.896 25.15 0.882
UEVD (Kim et al. 2022) F+E 29.93 0.905 24.08 0.869
REFID (Sun et al. 2023) F+E 30.10 0.913 24.55 0.875
Joint (Frame-based) LEDNet (Zhou et al. 2022) F 30.36 0.887 25.74 0.850
Joint (Event-guided) ELEDNet (Kim et al. 2024) F+E 31.30 0.925 25.04 0.873
@@
We evaluate CompEvent on the RELED real-world dataset (Kim et al. 2024) and the LOL-Blur synthetic dataset (Zhou et al. 2022). Training details and hyperparameters are provided in the supplementary material. (1) RELED (Kim et al. 2024) is the first large-scale real-world benchmark built for the joint low-light enhancement and motion deblurring tasks. The dataset is acquired through an optical beam splitting system that can simultaneously record low-light blurry videos, the corresponding high-quality clear images, and high-fidelity event streams. (2) LOL-Blur (Zhou et al. 2022) is a large-scale synthetic dataset that provides low-light blurry image and clear image pairs for the joint low-light enhancement and deblurring tasks. To adapt our event-based approach, we generate the corresponding event streams for it using the ESIM simulator (Rebecq et al. 2018).